40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 25 02:03:59 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts 
    Reply  Fri Jul 28 15:36:32 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts errSig.pdfQPD_simulink.pngfeedbackTopology.pdf
       Reply  Tue Aug 1 16:05:01 2017, gautam, Update, Optical Levers, Optical lever tuning - cost function construction loopPerformance.pdfOplevLoop.pngOL_errSigs.pdfDAC_saturation.pdf
          Reply  Thu Nov 16 13:57:01 2017, gautam, Update, Optical Levers, Optical lever noise VertexOLnoise.pdfITMYvsSRM.pdf
             Reply  Thu Nov 16 15:43:01 2017, rana, Update, Optical Levers, Optical lever noise 
                Reply  Tue Nov 21 11:37:29 2017, gautam, Update, Optical Levers, BS OL calibration updated OL_calib_BS_PERROR.pdfOL_calib_BS_YERROR.pdfVertexOLnoise_updated.pdf
                   Reply  Tue Nov 21 16:28:23 2017, gautam, Update, Optical Levers, BS OL calibration updated BS_oplev_sweep.png
                      Reply  Tue Nov 21 23:04:12 2017, gautam, Update, Optical Levers, Oplev "noise budget" vertexOL_noises.pdf
                         Reply  Wed Nov 22 05:41:32 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                            Reply  Wed Nov 22 15:29:23 2017, gautam, Update, Optical Levers, Oplev "noise budget" OplevTest.jpg
                               Reply  Wed Nov 22 16:40:00 2017, Koji, Update, Optical Levers, Oplev "noise budget" 
                               Reply  Wed Nov 22 19:20:01 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                                  Reply  Wed Nov 22 23:56:14 2017, gautam, Update, Optical Levers, Oplev "noise budget" 
                                     Reply  Thu Nov 23 18:03:52 2017, gautam, Update, Optical Levers, Oplev "noise budget" ITMY_P_noise.pdfITMY_Y_noise.pdf
Message ID: 13448     Entry time: Wed Nov 22 15:29:23 2017     In reply to: 13444     Reply to this: 13449   13451
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: Oplev "noise budget" 

[steve, gautam]

What is the best way to set this test up?

I think we need a QPD to monitor the spot rather than a single element PD, to answer this question about the sensor noise. Ideally, we want to shoot the HeNe beam straight at the QPD - but at the very least, we need a lens to size the beam down to the same size as we have for the return beam on the Oplevs. Then there is the power - Steve tells me we should expect ~2mW at the output of these HeNes. Assuming 100kohm transimpedance gain for each quadrant and Si responsivity of 0.4A/W at 632nm, this corresponds to 10V (ADC limit) for 250uW of power - so it would seem that we need to add some attenuating optics in the way.

Also, does anyone know of spare QPDs we can use for this test? We considered temporarily borrowing one of the vertex OL QPDs (mark out its current location on the optics table, and move it over to the SP table), but decided against it as the cabling arrangement would be too complicated. I'd like to use the same DAQ electronics to acquire the data from this test as that would give us the most direct estimate of the sensor noise for supposedly no motion of the spot, although by adding 3 optics between the HeNe and the QPD, we are introducing possible additional jitter couplings...

Quote:

For the OL NB, probably don't have to fudge any seismic noise, since that's a thing we want to suppress. More important is "what the noise would be if the suspended mirrors were no moving w.r.t. inertial space".

For that, we need to look at the data from the OL test setup that Steve is putting on the SP table.

 

Attachment 1: OplevTest.jpg  22 kB  | Hide | Hide all
OplevTest.jpg
ELOG V3.1.3-